### v1

# Nonlinear Stability of First-Order Relativistic Viscous Hydrodynamics

Nonlinear Stability of
First-Order Relativistic Viscous Hydrodynamics
Heinrich Freistühler   and   Matthias Sroczinski Department of Mathematics, University of Konstanz, 78457 Konstanz, Germany; Supported by DFG Grant No. FR 822/11-1
(May 31, 2024)
1 Introduction
In this paper we study the relativistic dynamics of a barotropic fluid under dissipation, as described by a system of partial differential equations
$\frac{\partial}{\partial x^{\beta}}\left(T^{\alpha\beta}+\Delta T^{\alpha\beta}\right)=0$ (1.1)
of second order. Here the ideal part of the stress-energy tensor is given by
$T^{\alpha\beta}=\theta p^{\prime}(\theta)u^{\alpha}u^{\beta}+p(\theta)g^{\alpha\beta},$ (1.2)
where $u^{\alpha}$ is the 4-velocity and the fluid is specified by an equation of state $p=p(\theta)$ that gives its pressure $p$ as a function of its temperature $\theta$, with
$p^{\prime}(\theta),p^{\prime\prime}(\theta)>0.$ (1.3)
Regarding the dissipative part, we will explore tensors of the general form [3, 4]
$-\Delta T^{\alpha\beta}\equiv u^{\alpha}u^{\beta}P+(\Pi^{\alpha\gamma}u^{\beta}+\Pi^{\beta\gamma}u^{\alpha})Q_{\gamma}+\Pi^{\alpha\beta}R+\Pi^{\alpha\gamma}\Pi^{\beta\delta}S_{\gamma\delta}$ (1.4)
where
$P=\tilde{\kappa}u^{\gamma}{\partial\theta\over\partial x^{\gamma}}+\tilde{\tau}{\partial u^{\gamma}\over\partial x^{\gamma}},\quad R=\tilde{\omega}u^{\gamma}{\partial\theta\over\partial x^{\gamma}}+\tilde{\chi}{\partial u^{\gamma}\over\partial x^{\gamma}},$
$Q_{\gamma}\equiv\tilde{\nu}{\partial\theta\over\partial x^{\gamma}}+\tilde{\mu}u^{\delta}{\partial u_{\gamma}\over\partial x^{\delta}},\quad S_{\gamma\delta}\equiv\tilde{\eta}\bigg{(}{\partial u_{\gamma}\over\partial x^{\delta}}+{\partial u_{\delta}\over\partial x^{\gamma}}-{2\over 3}g_{\gamma\delta}{\partial u^{\epsilon}\over\partial x^{\epsilon}}\bigg{)},$
$\Pi^{\alpha\beta}\equiv g^{\alpha\beta}+u^{\alpha}u^{\beta}$
and the dissipation coefficients
$\tilde{\kappa},\tilde{\tau},\tilde{\omega},\tilde{\chi},\tilde{\nu},\tilde{\mu},\tilde{\eta}$
are, in principle arbitrary, functions of $\theta$. Using the natural variable $\psi^{\alpha}=\theta^{-1}u^{\alpha}$ that ranges in $\mathcal{U}=\{\psi\in\mathbb{R}^{4}:\psi_{\alpha}\psi^{\alpha}<0\}$, we express (1.1) as
$A^{\alpha\beta\gamma}(\psi(x))\frac{\partial\psi_{\gamma}}{\partial x^{\beta}}(x)=\frac{\partial}{\partial x^{\beta}}\left(B^{\alpha\beta\gamma\delta}(\psi(x))\frac{\partial\psi_{\gamma}}{\partial x^{\delta}}\right),$ (1.5)
and consider solutions $\psi:[0,\infty)\times\mathbb{R}^{3}\to\mathcal{U}$ of (1.5) which satisfy initial conditions
$\psi(0,\cdot)={{}^{0}}\psi,\quad\frac{\partial\psi}{\partial x_{0}}(0,\cdot)={{}^{1}}\psi.$ (1.6)
with given data ${{}^{0}}\psi,{{}^{1}}\psi:\mathbb{R}^{3}\to\mathcal{U}$.
The purpose of the paper is to show nonlinear stability of homogeneous states in the following sense.
Theorem 1.
Fix a number $s>5/2$ and assume that normalized versions $(\kappa,\omega,\nu)=\theta^{-2}(\tilde{\kappa},\tilde{\omega},\tilde{\nu})$, $(\tau,\chi,\mu,\eta)=\theta^{-1}(\tilde{\tau},\tilde{\chi},\tilde{\mu},\tilde{\eta})$ of the dissipation coefficients satisfy
$\kappa,\mu,\eta,\nu\sigma>0\quad\text{ with }\quad\sigma=\chi-4\eta/3$
and either condition (C1), i.e,
$(\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu$ $>0,$ (C1.1)
$((\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu)^{2}-4\nu\mu\kappa\sigma$ $>0,$ (C1.2)
$c_{s}^{-2}(\omega+\tau)-\kappa-c_{s}^{-4}\sigma$ $>0,$ (C1.3)
$(\kappa+c_{s}^{-2}\mu)(\tau+\omega+\mu-c_{s}^{-2}\sigma)[(\tau+\mu)(\omega+\nu)-(\mu\nu+\kappa\sigma)]\qquad\qquad$
$-\kappa\mu(\tau+\omega+\mu-c_{s}^{-2}\sigma)^{2}-(\kappa+c_{s}^{-2}\mu)^{2}\nu\sigma$ $>0,$ (C1.4)
or condition (C2), i.e.,
$\kappa\sigma=\nu\mu$ $<0,\quad\tau+\mu=\omega+\nu=0,$ (C2.1)
$\sigma+c_{s}^{2}\mu<0,$ (C2.2)
where $c_{s}=\sqrt{p^{\prime}(\theta)/\theta p^{\prime\prime}(\theta)}>0$ is the speed of sound.
Then for any rest state $\bar{\psi}=(\bar{\theta}^{-1},0,0,0),\bar{\theta}>0$, there exist constants $\delta,C>0$ such that the following holds for all function pairs ${{}^{0}}\psi,{{}^{1}}\psi:\mathbb{R}^{3}\to\mathcal{U}$ with ${{}^{0}}\psi-\bar{\psi}\in H^{s+1}\cap L^{1}$, ${{}^{1}}\psi\in H^{s}\cap L^{1}$: If
$\|{{}^{0}}\psi-\bar{\psi}\|_{H^{s+1}},\|{{}^{1}}\psi\|_{H^{s}},\|{{}^{0}}\psi-\bar{\psi}\|_{L^{1}},\|{{}^{1}}\psi\|_{L^{1}}<\delta,$
then there exists a unique global solution $\psi$ of (1.5), (1.6) satisfying $\psi-\bar{\psi}\in C^{0}([0,\infty),H^{s+1})\cap C^{1}([0,\infty),H^{s})$ and, for all $t\in[0,\infty)$,
$\|\psi(t)-\bar{\psi}\|_{H^{s}}+\|\psi_{t}(t)\|_{H^{s-1}}$ $\leq C(1+t)^{-\frac{3}{4}}(\|{{}^{0}}\psi-\bar{\psi}\|_{H^{s}}+\|{{}^{1}}\psi\|_{H^{s-1}}+\|{{}^{0}}\psi-\bar{\psi}\|_{L^{1}}+\|{{}^{1}}\psi\|_{L^{1}}).$
I. e., at least for initial data that are sufficiently small perturbations of the homogeneous reference state $\bar{\psi}$, a unique solution to the nonlinear Cauchy problem exists globally in time and decays time-asymptotically to $\bar{\psi}$ at the rate $t^{-3/4}$.
In the statement of the theorem, the notations $C^{k},H^{l},$ and $L^{1}$ mean spaces of functions that are $k$ times continuously differentiable, square integrable together with their derivatives of up to $l$th order, or integrable, respectively, and .$\|.\|_{H^{l}},\|.\|_{L^{1}}$ are the natural norms in $H^{l}$ and $L^{1}$. For the argumentation in this paper we do not need to enter any details of these function spaces. We derive the assertion by appealing to a result of the second author [8] which reduces global existence and decay of solutions in these spaces for a general class of systems of the form (1.5) to two algebraic conditions, $(H)_{B}$ and $(D)$ below. Our concrete task here is thus the verification of these conditions under the assumptions (C1) or (C2).
Theorem 1 treats models introduced in [3, 4, 1] from a common perspective regarding specific aspects of their dissipativity. While the decay of Fourier modes as such had been considered in these papers, we here check the refined criterion (D) of [2] that, differently from the hyperbolic-parabolic situation of [6], is needed in the present hyperbolic-hyperbolic context in addition to mere decay of modes in order to ensure asymtotic stability in the sense of decay of solutions in appropriate function spaces.
For the models considered in [1], criterion (D) indeed again amounts to the same inequalities, namely (C.1), as are stated in that paper for the most relevant case $\nu=\mu$ [^1] Inequalities (C1.3), (C1.4) only look slightly different from (10a) and (10b) in [1] since some coefficients scale with $c_{s}^{2}$, thus $\chi_{1}=c_{s}^{2}\kappa$ and $\chi_{3}=c_{s}^{2}\omega$. But only after checking this, we have confirmed the decay in Sobolev spaces we report in Theorem 1.
In the following, we use the matrix notation
$B^{\beta\delta}=(B^{\alpha\beta\gamma\delta})_{0\leq\alpha,\gamma\leq 3},\quad A^{\beta}=(A^{\alpha\beta\gamma})_{0\leq\alpha,\gamma\leq 3}$
and the Fourier symbols
$B(\psi,\boldsymbol{\xi})$ $=B^{jk}\xi_{j}\xi_{k},\quad C(\psi,\boldsymbol{\xi})=(B^{0j}(\psi)+B^{j0}(\psi))\xi_{j},$
$A(\psi,\boldsymbol{\xi})$ $=A^{j}(\psi)\xi_{j},\leavevmode\nobreak\ \leavevmode\nobreak\ \boldsymbol{\xi}=(\xi_{1},\xi_{2},\xi_{3})\in\mathbb{R}^{3},\leavevmode\nobreak\ \xi=|\boldsymbol{\xi}|.$
2 Hyperbolicity
By a classical result of Taylor [9] the Cauchy problem (1.5), (1.6) is locally well-posed if the differential operator
$\psi\mapsto\frac{\partial}{\partial x^{\beta}}\left(B^{\alpha\beta\gamma\delta}(\psi(x))\frac{\partial\psi_{\gamma}}{\partial x^{\delta}}\right)$ (2.1)
is (second-order) hyperbolic in the following sense:
  1. (HB)
    (a) $B^{00}$ is negative definite,
    (b) with $\check{B}^{\beta\delta}$ defined through
    $\check{B}^{\beta\delta}=(-B^{00})^{-1/2}B^{\beta\delta}(-B^{00})^{-1/2},$
    the matrix family
    $i\mathcal{B}(\psi,\boldsymbol{\omega})=i\begin{pmatrix}0&I_{4}\\ \check{B}(\psi,\boldsymbol{\omega})&i\check{C}(\psi,\boldsymbol{\omega})\end{pmatrix},\quad(\psi,\boldsymbol{\omega})\in\mathcal{U}\times S^{2},$
    permits a symbolic symmetrizer $\mathcal{S}$.
Recall that for $K\in C^{\infty}(\mathcal{U}\times S^{d-1},\mathbb{C}^{n\times n})$, a symbolic symmetrizer is a smooth mapping $S\in C^{\infty}(\mathcal{U}\times S^{d-1},\mathbb{C}^{n\times n})$, bounded as well as all its derivatives, such that for some $c>0$ and all $(\psi,\boldsymbol{\omega})\in\mathcal{U}\times S^{d-1}$
$S(\psi,\boldsymbol{\omega})=S(\psi,\boldsymbol{\omega})^{*}\geq cI_{n},\leavevmode\nobreak\ \leavevmode\nobreak\ S(\psi,\boldsymbol{\omega})K(\psi,\boldsymbol{\omega})=(S(\psi,\boldsymbol{\omega})K(\psi,\omega))^{*}.$
Clearly, if a matrix family $K$ admits a symbolic symmetrizer all eigenvalues of $K(\psi,\boldsymbol{\omega})$ are real and semi-simple. On the other hand a matrix family $K$ admits a symbolic symmetrizer if the latter holds and additionally the multiplicity of the eigenvalues of $K(\psi,\boldsymbol{\omega})$ is constant on $\mathcal{U}\times S^{d-1}$. This motivates the following notion.
Definition. We call $B^{\alpha\beta\gamma\delta}$ or, more precisely, the differential operator (2.1), semi-strictly hyperbolic if (a) holds and for all $(\psi,\boldsymbol{\omega})\in\mathcal{U}\times S^{2}$ the eigenvalues of $i\mathcal{B}(\psi,\boldsymbol{\omega})$ are real semi-simple with multiplicities independent of $(\psi,\boldsymbol{\omega})$.
Since we will show that $B^{\alpha\beta\gamma\delta}$ is hyperbolic if and only if it is semi-strictly hyperbolic, it is sufficient to show (HB) in the rest frame due to Lorentz invariance. In the following all matrices are evaluated at $\bar{\psi}$ without further indication.
First note that the rest-frame representations are explicitly given as
$B^{00}=-\begin{pmatrix}\kappa&0\\ 0&\mu I_{3}\end{pmatrix}$
and
$B(\boldsymbol{\xi})=-\begin{pmatrix}\nu\xi^{2}&0\\ 0&(\chi-\frac{1}{3}\eta)\xi\xi^{t}-\eta\xi^{2}I_{3}\end{pmatrix},\quad C(\boldsymbol{\xi})=-\begin{pmatrix}0&(\tau+\mu)\boldsymbol{\xi}^{t}\\ (\omega+\nu)\boldsymbol{\xi}&0\end{pmatrix}.$
It is straightforward to see that for any $\boldsymbol{\xi}\in\mathbb{R}^{3}$ the matrices $B^{00},B(\boldsymbol{\xi}),C(\boldsymbol{\xi})$ all decompose in the sense of linear operators as $B^{00}=B^{00}_{\parallel}\oplus B^{00}_{\perp}$, $B(\boldsymbol{\xi})=B_{\parallel}(\xi)\otimes B_{\perp}(\xi)$, $C(\xi)=C_{\parallel}(\xi)\oplus C_{\perp}(\xi)$ with respect to the orthogonal decomposition $\mathbb{C}^{4}=(\mathbb{C}\times\mathbb{C}\boldsymbol{\xi})\oplus(\{0\}\times\{\boldsymbol{\xi}\}^{\perp})$, where
$B^{00}_{\parallel}$ $=-\begin{pmatrix}\kappa&0\\ 0&\mu\end{pmatrix},$ $B^{00}_{\perp}$ $=-\mu I_{2},$
$B_{\parallel}(\xi)$ $=-\xi^{2}\begin{pmatrix}\nu&0\\ 0&\sigma\end{pmatrix},$ $B_{\perp}(\xi)$ $=\xi^{2}\eta I_{2},$
$C_{\parallel}(\xi)$ $=-\xi\begin{pmatrix}0&\tau+\mu\\ \omega+\nu&0\end{pmatrix},$ $C_{\perp}(\xi)$ $=0,$
so that we can treat dissipativity on ×$\mathbb{C}\times\mathbb{C}\boldsymbol{\xi}$ and $\{0\}\times\{\boldsymbol{\xi}\}^{\perp}$ separately. We also just write $B_{\parallel}$ instead of $B_{\parallel}(1)$, etc., if there is no concern for confusion. .
Lemma 1.
$B^{\alpha\beta\gamma\delta}$ satisfies (HB) if and only if $\kappa,\mu,\eta>0$ and either
  • (i)
    $\nu\sigma>0$ and (C1.1) and (C1.2), or
  • (ii)
    $\nu\sigma>0$ and (C2.1), or
  • (iii)
    $\nu=\sigma=0\leavevmode\nobreak\ \text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }(\tau+\mu)(\omega+\nu)>0$
are true.
Proof.
Trivially, $-B^{00}<0$ is equivalent to $\kappa,\mu>0$; we assume this to be the case. Next, we treat (HB) (b), separately on the spaces ×$\mathbb{C}\times\mathbb{C}\boldsymbol{\xi}$ and $\{0\}\times\{\boldsymbol{\xi}\}^{\perp}$. We first find that
$\mathcal{B}_{\perp}=\begin{pmatrix}0&I_{2}\\ -\bar{B}_{\perp}&i\bar{C}_{\perp}\end{pmatrix}=\begin{pmatrix}0&I_{2}\\ -(\eta/\mu)I_{2}&0\end{pmatrix}$
has the eigenvalues ±$\pm\sqrt{-\eta/\mu}$, which are purely imaginary and semi-simple (of multiplicity two) if and only if $\eta>0$.
Next, we see that $\lambda=ib$, $b\in\mathbb{R}$, is an eigenvalue of
$\mathcal{B}_{\parallel}=\begin{pmatrix}0&I_{2}\\ -\bar{B}_{\parallel}&i\bar{C}_{\parallel}\end{pmatrix}$
if and only if
$0=\det(-B_{\parallel}^{00}b^{2}-B_{\parallel}-bC_{\parallel})=\kappa\mu b^{4}-((\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu)b^{2}+\nu\sigma:=\mathfrak{p}(b^{2}).$ (2.2)
Clearly, the quadratic $\mathfrak{p}$ has only positive real roots, i.e. $\mathcal{B}_{\parallel}$ has only purely imaginary eigenvalues, if and only if
$(\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu\geq 0,$
$((\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu)^{2}-4\nu\mu\kappa\sigma\geq 0.$
If $\nu\sigma>0$, (C1.1) and (C1.2) hold, these eigenvalues are all distinct. Next, note that for an eigenvalue $\lambda=ib$ of $\mathcal{B}_{\parallel}$ the eigenvectors are of the form $(V,ibV)$ with
$V\in\ker(B_{\parallel}^{00}b^{2}+B_{\parallel}+bC_{\parallel}).$
Thus $\lambda$ is semi-simple of multiplicity $2$ if and only if
$B_{\parallel}^{00}b^{2}+B_{\parallel}+bC_{\parallel}=0.$ (2.3)
If $\lambda=0$ is a semi-simple eigenvalue of $\mathcal{B}_{\parallel}$ this implies $\nu=\sigma=0$. And in this case the other eigenvalues are purely imaginary and semi-simple if and only if $(\tau+\mu)(\omega+\nu)>0$. Lastly, due to (2.3) $\mathcal{B}_{\parallel}$ has two non-zero semi-simple eigenvalues of multiplicity $2$ if and only if (C2.1) holds. ∎
As mentioned the lemma shows that $B^{\alpha\beta\gamma\delta}$ is hyperbolic if and only if it is semi-strictly hyperbolic. In this case $\mathcal{B}(\omega)$ always has two eigenvalues of multiplicity $2$, which in the rest frame correspond to the eigenvalues of $\mathcal{B}_{\perp}$, and (i), (ii), (iii) above correspond to different multiplicities of the eigenvalues of $\mathcal{B}_{\parallel}$. In (i) $\mathcal{B}_{\parallel}$ all eigenvalues have multiplicity 1, in (ii) there are two distinct eigenvalues with multiplicity 2 and in case (iii) 00 is an eigenvalue of multiplicity $2$ and there exist two distinct non-zero eigenvalues of multiplicity $1$. As (iii) implies $\mathcal{B}_{\parallel}=0$ this is not physical and we will not treat this case any further.
Both cases (i) and (ii) have been studied before. Situation (i) with the further assumption $\nu=\mu$ recovers the equations proposed and investigated in [1]. In situation (ii) (2.1) is second order symmetric hyperbolic in the sense defined in [5]; this case comprises the equations proposed in [4], where additionally $\mu=\sigma$ [^2] Note that what is $-\sigma$ here is called $\sigma$ in [4].
3 Dissipativity
It is well-known and also easily checked that under conditions (1.3) the relativistic Euler operator
$\psi\mapsto A^{\alpha\beta\gamma}(\psi)\frac{\partial\psi_{\gamma}}{\partial x^{\delta}}$
is (first-order) symmetric hyperbolic, i.e., the matrices $A^{\beta}$ are symmetric with $A^{0}$ positive definite. Thus if $B^{\alpha\beta\gamma\delta}$ satisfies (HB) system (1.5) is of hyperbolic-hyperbolic type as introduced in [2] and the non-linear stability of the rest state $\bar{\psi}=\bar{\theta}^{-1}(1,0,0,0)^{t}$, $\bar{\theta}>0$ fixed, is characterized by the following condition (we use $\bar{B}^{\alpha\beta\gamma\delta}=\check{B}^{\alpha\beta\gamma\delta}(\bar{\psi})$ etc.)
Condition (D). The tensors $\bar{B}^{\alpha\beta\gamma\delta}$, $\bar{A}^{\alpha\beta\gamma}$ satisfy three conditions:
  •  (D1)
    For every $\boldsymbol{\omega}\in S^{2}$, all restrictions, as a quadratic form, of
    $W_{1}=(\bar{A}^{0})^{-1/2}\big{(}-\bar{B}(\boldsymbol{\omega})+(\bar{A}^{0})^{-1}\bar{A}(\boldsymbol{\omega})(\bar{A}^{0})^{-1}\bar{A}(\boldsymbol{\omega})+\bar{C}(\boldsymbol{\omega})(\bar{A}^{0})^{-1}\bar{A}(\boldsymbol{\omega})\big{)}(\bar{A}^{0})^{-1/2}$
    on the eigenspaces $E=J_{E}^{-1}(\mathbb{C}^{n})$ of
    $W_{0}=-(\bar{A}^{0})^{-1/2}\bar{A}(\boldsymbol{\omega})(\bar{A}^{0})^{-1/2}$
    are uniformly negative in the sense that
    $J_{E}^{*}\left(W_{1}+W_{1}^{*}\right)J_{E}\leq-\bar{c}\ J_{E}^{*}J_{E}\quad\text{with one }\bar{c}>0.$
  • (D2)
    There exists a symbolic symmetrizer $\mathcal{S}$ for $i\mathcal{B}$ such that for every $\boldsymbol{\omega}\in S^{2}$, all restrictions, as a quadratic form, of
    $\mathcal{W}_{1}=\bar{\mathcal{S}}(\boldsymbol{\omega})^{\frac{1}{2}}\begin{pmatrix}0&0\\ -i\bar{A}(\boldsymbol{\omega})&-\bar{A}^{0}\end{pmatrix}\bar{\mathcal{S}}(\boldsymbol{\omega})^{-\frac{1}{2}}$
    on the eigenspaces $\mathcal{E}=\mathcal{J}_{\mathcal{E}}^{-1}(\mathbb{C}^{2n})$ of
    $\mathcal{W}_{0}=\bar{\mathcal{S}}(\boldsymbol{\omega})^{\frac{1}{2}}\bar{\mathcal{B}}(\boldsymbol{\omega})\bar{\mathcal{S}}(\boldsymbol{\omega})^{-\frac{1}{2}}$
    are uniformly negative in the sense that
    $\mathcal{J}_{\mathcal{E}}^{*}\left(\mathcal{W}_{1}+\mathcal{W}_{1}^{*}\right)\mathcal{J}_{\mathcal{E}}\leq-\bar{c}\ \mathcal{J}_{\mathcal{E}}^{*}\mathcal{J}_{\mathcal{E}}\quad\text{with one }\bar{c}>0.$
  • (D3)
    All solutions $(\lambda,\boldsymbol{\xi})\in\mathbb{C}\times(\mathbb{R}^{d}\setminus\{0\})$ of the dispersion relation for (1.5) have Re$(\lambda)<0$.
We note that with respect to the orthogonal decomposition $\mathbb{C}^{4}=(\mathbb{C}\times\boldsymbol{\xi}\mathbb{C})\oplus(\{0\}\times\{\boldsymbol{\xi}\}^{\perp})$ also $A^{0}$ and $A(\boldsymbol{\xi})$ decompose in the sense of linear operators as
$A^{0}=A^{0}_{\parallel}\oplus A^{0}_{\perp},\qquad A(\boldsymbol{\xi})=A_{\parallel}(\xi)\oplus A_{\perp}(\xi).$
So we can treat also the dissipativity on $\{0\}\times\{\boldsymbol{\xi}\}^{\perp}$ and ×$\mathbb{C}\times\mathbb{C}\boldsymbol{\xi}$ separately.
For the case mentioned in the last sentence of the previous section, non-linear stability of the rest state was first shown in [7] – a generalization for all systems that satisfy (ii) is straightforward:
Lemma 2.
Assume $\kappa,\mu,\eta>0$ and (C2.1). Then (D1), (D2) and (D3) are all equivalent to (C2.2).
Proof.
By the arguments presented in [2], Section 3.2, (D) holds in general for $-B^{00},B(\boldsymbol{\omega}),A^{0}>0$ and $A(\boldsymbol{\omega})=C(\boldsymbol{\omega})=0$. Thus it holds here on $\{0\}\times\{\boldsymbol{\xi}\}^{\perp}$.
To study it on ×$\mathbb{C}\times\mathbb{C}\boldsymbol{\xi}$, set $\tilde{\sigma}:=-\sigma/\mu$ and note that (C2.1)implies $B_{\parallel}=\tilde{\sigma}B_{\parallel}^{00}$, i.e., $\bar{B}_{\parallel}=\tilde{\sigma}I_{2}$. As $\tilde{\sigma}>0$ conditions (D1), (D2), (D3) are all equivalent to
$\tilde{\sigma}^{\frac{1}{2}}\bar{A}^{0}_{\parallel}\pm\bar{A}_{\parallel}(\xi)>0$
(cf. [2], Section 3.2), which directly yields the assertion. ∎
In the following, we use the abbreviation $r=c_{s}^{-2}$.
Lemma 3.
Assume $\kappa,\mu,\eta>0$, $\nu\sigma>0$ and (C1.1), (C1.2). Then the following hold
  1. (i)
    (D3) is equivalent to
    $r(\omega+\tau)-\kappa-r^{2}\sigma\geq 0,$ (3.4)
    $(\kappa+r\mu)(\tau+\omega+\mu-r\sigma)[(\tau+\mu)(\omega+\nu)-(\mu\nu+\kappa\sigma)]$ (3.5)
    $-\kappa\mu(\tau+\omega+\mu-\hat{r}\eta)^{2}-(\kappa+r\mu)^{2}\nu\sigma$ $\geq 0.$
    where at least one of the inequalities is strict.
  2. (ii)
    (D1) is equivalent to (C1.3).
  3. (iii)
    Assume (D3). Then (D2) is equivalent to
    $(\kappa+r\mu)(\tau+\omega+\mu-r\sigma)[(\tau+\mu)(\omega+\nu)-(\mu\nu+\kappa\sigma)]$ (3.6)
    $-\kappa\mu(\tau+\omega+\mu-\hat{r}\eta)^{2}-(\kappa+r\mu)^{2}\nu\sigma$ $\neq 0.$
Proof.
As mentioned in the last proof, (D1), (D2), (D3) are satisfied on $\{0\}\times\{\boldsymbol{\xi}\}^{\perp}$ under the assumptions $\mu,\eta>0$. Thus we only need to consider the symbols on ×$\mathbb{C}\times\mathbb{C}\boldsymbol{\xi}$; we drop the subscript $\parallel$.
(i) The dispersion relation is given as
00 0 $=\det(-\lambda^{2}B^{00}+B(\xi)-i\lambda C(\xi)+\lambda A^{0}+iA(\xi)$
$=\kappa\mu\lambda^{4}+(\kappa+r\mu)\lambda^{3}+(\xi^{2}((\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu)+r)\lambda^{2}+\lambda\xi^{2}(\tau+\omega+\mu-r\sigma)$
$\quad+\xi^{4}\nu\sigma+\xi^{2}.$
We directly see that by rescaling $\kappa$ by $r^{2}$, $\mu,\nu,\omega,\tau$ by $r$ and $\xi^{2},\lambda$ by $r^{-1}$ the dispersion relation and (C1.3), (C1.4) become independent of $r$. Thus we may assume $r=1$ w.l.o.g.. We use $\alpha=\xi^{2}$ and set
$a_{0}$ $=\alpha^{2}\nu\sigma+\alpha,\leavevmode\nobreak\ \leavevmode\nobreak\ a_{1}=\alpha(\tau+\omega+\mu-\sigma),$
$a_{2}$ $=\alpha((\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu)+1,\leavevmode\nobreak\ \leavevmode\nobreak\ a_{3}=\kappa+\mu,\leavevmode\nobreak\ \leavevmode\nobreak\ a_{4}=\kappa\mu.$
By the Routh-Hurwitz criterion, (D3) is equivalent to $a_{i}>0$, $i=0,\ldots 4$, and
$\Delta(a):=a_{1}a_{2}a_{3}-(a_{1}^{2}a_{4}+a_{3}^{2}a_{0})>0$
for all $\alpha\in(0,\infty)$. Clearly $a_{0},a_{3},a_{4}>0$ by assumption and $a_{2}>0$ by (C1.1). Furthermore
$\Delta=a(\Delta^{(1)}+a\Delta^{2}{(2)}),$
where
$\Delta^{(1)}$ $=\omega+\tau-\kappa-\sigma$
$\Delta^{(2)}$ $=(\kappa+\mu)(\tau+\omega+\mu-\eta)[(\tau+\mu)(\omega+\nu)-(\mu\nu+\kappa\sigma)]-\kappa\mu(\tau+\omega+\mu-\sigma)^{2}-(\kappa+\mu)^{2}\nu\sigma.$
As $\Delta^{(1)}\geq 0$ implies $a_{1}>0$ this finishes the proof of (i).
(ii) To show (ii) and (iii) we perform the transformation $M\to(-B^{00})^{-\frac{1}{2}}M(-B^{00})^{-\frac{1}{2}}$ on $B^{00},B,C,A^{0},A$ to obtain (without changing the notation)
$B^{00}=-I_{2},B=B(1)=-\begin{pmatrix}\frac{\nu}{\kappa}&0\\ 0&\frac{\sigma}{\mu}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ C=C(1)=-\frac{1}{\sqrt{\kappa\mu}}\begin{pmatrix}0&\tau+\nu\\ \omega+\mu\end{pmatrix},$
$A^{0}=\begin{pmatrix}\frac{1}{\kappa}&0\\ 0&\frac{1}{\mu}\end{pmatrix},\quad A=A(1)=\frac{1}{\sqrt{\kappa\mu}}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$
The eigenspaces of
$W_{0}=(A^{0})^{-\frac{1}{2}}A(A^{0})^{-\frac{1}{2}}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$
are $E_{\pm}=\mathbb{C}(\pm 1,1)^{t}$ and the restriction of
$W_{1}=(A^{0})^{-\frac{1}{2}}(-B+(A^{0})^{-1}A(A^{0})^{-1}A+C(A^{0})^{-1}A)(A^{0})^{-\frac{1}{2}}$
on $E_{\pm}$ is
$\sigma+\kappa-(\omega+\tau),$
which proves (ii).
(iii) For better readability we only give the proof for $(\tau+\mu)(\omega+\nu)\neq 0$ at this point. The other case follows with analogous arguments. By (HB) the matrix
$\mathcal{B}=\begin{pmatrix}0&I\\ -B&iC\end{pmatrix}$
has four simple purely imaginary eigenvalues, which are given by
$\lambda_{\pm s}=\pm i\gamma_{s}=\pm\sqrt{-\beta_{s}},\leavevmode\nobreak\ \leavevmode\nobreak\ s=1,2$
where $0<\beta_{1}<\beta_{2}$ are the roots of the polynomial $\pi$ with $\pi$ being defined in (2.2). To determine the left and right eigenvectors write
$\mathcal{B}=\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ b&0&0&-ia\\ 0&c&-id&0\end{pmatrix},$
where
$a=\frac{\tau+\mu}{\sqrt{\kappa\mu}},\leavevmode\nobreak\ \leavevmode\nobreak\ b=\frac{\nu}{\kappa},\leavevmode\nobreak\ \leavevmode\nobreak\ c=\frac{\sigma}{\mu},\leavevmode\nobreak\ \leavevmode\nobreak\ d=\frac{\omega+\nu}{\sqrt{\kappa\mu}}.$
Since $(\tau+\mu)(\omega+\nu)\neq 0$ we find that for $s=-2,-1,1,2$ a left respectively right eigenvector corresponding to $\lambda_{s}$ is given as
$L_{s}=\begin{pmatrix}-db\gamma_{s}&(da-b-\gamma_{s}^{2})\gamma_{s}^{2}&-id&i(\gamma_{s}^{2}+b)\gamma_{s}\end{pmatrix},\quad R_{s}=\begin{pmatrix}a\gamma_{s}\\ -(b+\gamma_{s}^{2})\\ ia\gamma_{s}^{2}\\ -i(b+\gamma_{s}^{2})\gamma_{s}\end{pmatrix}.$
Assumiing (D3) condition (D2) is now equivalent to
$L_{s}\mathcal{A}R_{s}\neq 0,s=-2,-1,1,2.$
We find for $s=1,2$
$L_{\pm s}\mathcal{A}R_{\pm_{s}}=\frac{\beta_{s}}{\mu}\left[-\beta_{s}^{2}+\frac{1}{\kappa}\left(\tau+\omega+\mu-\nu-\frac{1}{\kappa}(\tau+\nu)(\omega+\mu)\right)\beta_{s}+\frac{(\tau+\omega+\mu)\nu}{\kappa^{2}}\right]=:\frac{\beta_{s}}{\mu}q(\beta_{s}).$
With
$K:=\frac{(\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu}{\kappa\mu},\quad L:=\frac{\nu\sigma}{\kappa\mu}$ (3.7)
the zeros of the polynomial $\pi$ are given as
$\beta_{1}$ $=\frac{1}{2}\left(K-\sqrt{K^{2}-4L}\right),\quad\beta_{2}=\frac{1}{2}\left(K+\sqrt{K^{2}-4L}\right).$
Also using
$M=\frac{1}{\kappa}\left(\tau+\omega+\mu-\nu-\frac{1}{\kappa}(\tau+\nu)(\omega+\mu)\right),\quad N=\frac{(\tau+\omega+\mu)\nu}{\kappa^{2}}$
we get that $q(\beta_{s})\neq 0$ is equivalent to
±$\pm(2(L+N)+K(M-K))\neq(K-M)\sqrt{K^{2}-4L}.$
Squaring both sides gives
$(L+N)^{2}+(M-K)(ML+KN)\neq 0$
and straightforward calculations give
$\frac{\kappa^{5}\mu^{2}}{\nu(\tau+\nu)(\omega+\mu)}((L+N)^{2}+(M-K)(ML+KN))$
$=(\kappa+\mu)(\tau+\omega+\mu-\eta)[(\tau+\mu)(\omega+\nu)-(\mu\nu+\kappa\sigma)]-\kappa\mu(\tau+\omega+\mu-\sigma)^{2}-(\kappa+\mu)^{2}\nu\sigma,$
which shows the assertion. ∎
4 Causality
Equations describing the dynamics of fluids in the relativistic regime need to be causal, i.e. information must not propagate faster than the speed of light. This did not play a role in the previous argumentation as hyperbolicity and dissipation can be achieved independently of causality. However, at this point we will also formulate conditions equivalent to causality in order to characterize the regime of parameters for which the equations are physically relevant.
Definition 1.
The hyperbolic differential operator
$\psi\mapsto\frac{\partial}{\partial x^{\beta}}\left(B^{\alpha\beta\gamma\delta}(\psi)\frac{\partial\psi_{\gamma}}{\partial x^{\delta}}\right)$
is causal if at any $\psi\in\mathcal{U}$ all solutions $(\varphi_{0},\varphi_{1},\varphi_{2},\varphi_{3})\in\mathbb{R}^{4}\setminus\{0\}$ of
$\det(B^{\alpha\beta\gamma\delta}(\psi)\varphi_{\beta}\varphi_{\delta})=0$
are spacelike.
Note that causality is a Lorentz invariant property and thus again we only need to study the rest-frame representation of $B^{\alpha\beta\gamma\delta}$.
Lemma 4.
Let the operator (2.1) be hyperbolic. Then it is causal if and only if $\eta/\mu\leq 1$ and
$(\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu\leq\kappa\mu+\min\{\kappa\mu,\nu\sigma\}.$ (4.1)
Proof.
It is obvious that
$\det(B^{\alpha\beta\gamma\delta}(\bar{\psi})\varphi_{\beta}\varphi_{\delta})=0$
is equivalent to the fact that $\varphi_{0}$ is a characteristic speed of $i\mathcal{B}(\bar{\psi},\varphi_{1},\varphi_{2},\varphi_{3})$ As seen in the proof of Lemma 1 these speeds are, in the rest frame,
$\lambda^{0}_{\pm}=\pm\sqrt{\frac{\eta}{\mu}},\quad\lambda_{\pm}^{j}=\pm\sqrt{\beta}_{j},\leavevmode\nobreak\ \leavevmode\nobreak\ j=1,2$
with
$\beta_{1}$ $=\frac{1}{2}\left(K-\sqrt{K^{2}-4L}\right),\quad\beta_{2}=\frac{1}{2}\left(K+\sqrt{K^{2}-4L}\right)$
$K$ $=\frac{(\tau+\mu)(\nu+\omega)-\kappa\sigma-\nu\mu}{\kappa\mu},\quad L=\frac{\nu\sigma}{\kappa\mu}.$
Trivially, $|\lambda^{0}|\leq 1$ if and only if $\eta\leq\mu$, and straightforward calculations show that the property $|\beta^{j}|\leq 1$, $j=1,2$, is equivalent to (4.1).